OpenStudy (anonymous):

integral 9+6x/1+x^2 *dx evaluate the integral

OpenStudy (lgbasallote):

interesting..

OpenStudy (turingtest):

\[\int{9+6x\over1+x^2}dx=3\int{3+2x\over1+x^2}dx=3\int{3\over1+x^2}+ {2x\over1+x^2}dx\]that should give you a few ideas

OpenStudy (lgbasallote):

you can also do \[\int \frac{9+6x}{1+x^2}dx\implies \int \frac{9}{1+x^2}dx + \int \frac{6x}{1+x^2}dx\] same thing really

OpenStudy (anonymous):

I make u=1+x^2

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

go on..

OpenStudy (anonymous):

du=2x but i don"t really know what to do with the top

OpenStudy (turingtest):

look at both lgb's way and mine try to see how they are the same, and how you can reconcile your problem by comparing the two

OpenStudy (lgbasallote):

hint: you can only use u-sub in the second integral. the first integral is trig sub